OpenStudy (migitmack):

What is the slope of a line that passes through the point (−1, 1) and is parallel to a line that passes through (4, 6) and (−1, −4)?

OpenStudy (anonymous):

Find the slope of parallel line first passing through (4, 6) and (-1, -4)..

OpenStudy (anonymous):

\[\large Slope = \frac{y_2 - y_1}{x_2 - x_1}\]

OpenStudy (migitmack):

Slope 4/-5 ?

OpenStudy (migitmack):

cuz 6-4/-1-4

OpenStudy (anonymous):

Check for numerator...

OpenStudy (migitmack):

what

OpenStudy (anonymous):

Here \[\large y_2 = -4, y_1 = 6, x_2 = -1 , x_1 = 4\] Now put the values...

OpenStudy (migitmack):

woops its not 4 its 2

OpenStudy (migitmack):

2/-5

OpenStudy (anonymous):

\[Slope = \frac{-4 - 6}{-1-4} = ??\]

OpenStudy (anonymous):

It's y - y over x - x. -4 - 6 = ? over -1 - 4 = ?

OpenStudy (migitmack):